# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v1

@@ Introduction
A2. The Proof of Theorem 1
Theorem 1.
@@
A3. Classifiers Estimation Derivations
@@
Transition Classifier Derivation.
@@
Action Classifier Derivation.
@@
The action classifier is defined as $C_{\psi}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
$C_{\psi}(s,a)=P(\mathbb{A}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{A})P(\mathbb{A}|s_{t})P(s_{t})}{P(s_{t},a_{t})},$ (16)
$1-C_{\psi}(s,a)=P(\mathbb{C}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{C})P(\mathbb{C}|s_{t})P(s_{t})}{P(s_{t},a_{t})}.$ (17)
@@
$\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}=\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}.$ (18)
Similarly, $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}$ can be also approximated by the constant $\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|}$. According to the sampling method in Algorithm 2, we have
$\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}=\frac{d(s_{t})\left|\mathcal{D}_{\pi}\right|}{d(s_{t})\left|\mathcal{D}_{env}\right|}=\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|},$
where $d(s_{t})$ denotes the distribution of states in the offline dataset, and $\mathcal{D}_{\pi}$ represents the dataset collected by policy $\pi$. Thus, taking the logarithm of both sides of (18), we get
$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}-c_{2},$
where $c_{2}$ is a constant that does not influence the training process.
Appendix B B. Implementation Details
@@
Our model training approach is consistent with the methodology used in prior works (Janner et al. 2019; Yu et al. 2020). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. During model rollouts, we randomly choose one model from the best five models.
B2. Classifiers Training
@@
$\mathcal{L}_{\psi}=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi}(s,a)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi}(s,a))\right].$ (20)
@@
B3. Policy Learning
The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al. 2023). For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$. The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 2.
@@
Table 2: Hyperparameters for policy optimization in SAMBO
Appendix C C. Experimental Details
C1. Benchmark
We conduct experiments on Gym tasks (“v2” version) within the D4RL benchmark and now provide the sources for the reported performance on this benchmark. For TD3+BC (Fujimoto and Gu 2021), EDAC (An et al. 2021), TT (Janner, Li, and Levine 2021), RAMBO (Rigter, Lacerda, and Hawes 2022), and MOBILE (Sun et al. 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym. Since COMBO (Yu et al. 2021) does not provide source code, we include the results as reported in the original paper. The results of CQL (Kumar et al. 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym. The results of MOPO* are obtained from experiments using the OfflineRL-Kit libraryhttps://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets. These scores are referenced in Table 1.
C2. Hyperparameters
We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 3.
@@
walker2d-random 0.03 0.03 5 0
halfcheetah-medium 0.01 0.01 5 0
@@
hopper-medium-replay 0.02 0.01 5 0
walker2d-medium-replay 0.01 0.01 1 -0.2
halfcheetah-medium-expert 0.04 0.01 5 0
hopper-medium-expert 0.04 0.02 5 0
walker2d-medium-expert 0.05 0.05 1 0
Table 3: Hyperparameters of SAMBO used in the D4RL datasets.
Model Bias Coefficient $\alpha$.
We tune $\alpha$ in the range of $[0.01,0.05]$. For training stability, in the walker2d-medium dataset, we fix the transition classifier after 1 million iterations by setting $\alpha=0$. In the walker2d-medium-expert dataset, we apply an annealing factor to the transition classifier. Initially set to 3, the factor remains constant for the first 1 million iterations and then decays reciprocally to 1 over the subsequent 2 million iterations.
Policy Shift Coefficient $\beta$.
We tune $\beta$ in the range of $[0.01,0.05]$.
@@
C3. Computing Infrastructure

### v2

@@ A.1 The Supplement Analysis of Distribution Shift
A.2 The Proof of Theorem 1
Theorem A.1.
@@ A.2 The Proof of Theorem 1
A.3 Classifiers Estimation Derivations
@@
A.3.1 Transition Classifier Derivation.
@@
A.3.2 Action Classifier Derivation.
@@
The action classifier is defined as $C_{\psi_{1}}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
$C_{\psi_{1}}(s,a)=P(\mathbb{A}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{A})P(\mathbb{A}|s_{t})P(s_{t})}{P(s_{t},a_{t})},$ (16)
$1-C_{\psi_{1}}(s,a)=P(\mathbb{C}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{C})P(\mathbb{C}|s_{t})P(s_{t})}{P(s_{t},a_{t})}.$ (17)
@@
$\frac{C_{\psi_{1}}(s,a)}{1-C_{\psi_{1}}(s,a)}=\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}.$ (18)
Similarly, $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}$ can also be approximated by a classifier $C_{\psi_{2}}(s)$, i.e. $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}=\frac{C_{\psi_{2}}(s)}{1-C_{\psi_{2}}(s)}$. Thus, we can approximate $\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}$ using two classifiers,
$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi_{1}}(s,a)}{1-C_{\psi_{1}}(s,a)}-\log\frac{C_{\psi_{2}}(s)}{1-C_{\psi_{2}}(s)}.$
For simplicity, we use the notation
$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
Appendix B Implementation Details
@@ B.1 Model Training
Our model training approach is consistent with the methodology used in prior works (Janner et al., 2019; Yu et al., 2020). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. During model rollouts, we randomly choose one model from the best five models.
B.2 Classifiers Training
@@
[…]
@@
B.3 Policy Learning
The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al., 2023). For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$. The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 3.
Table 3: Hyperparameters for policy optimization in SAMBO
@@ B.3 Policy Learning
Appendix C Experimental Details
C.1 Benchmark
We conduct experiments on Gym tasks (“v2” version) within the D4RL (Fu et al., 2020) benchmark. We also evaluated SAMBO on the NeoRL (Qin et al., 2022) benchmark, a more challenging dataset generated using a more conservative policy to better reflect real-world data collection scenarios. These datasets are characterized by their narrow scope and limited coverage.
D4RL.
We now provide the sources for the reported performance on this benchmark. For TD3+BC (Fujimoto & Gu, 2021), EDAC (An et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym. Since COMBO (Yu et al., 2021) does not provide source code, we include the results as reported in the original paper. The results of CQL (Kumar et al., 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym. The results of MOPO* are obtained from experiments using the OfflineRL-Kit library [^2] https://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets. These scores are referenced in Table 1.
NeoRL.
We report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores reported in the MOBILE (Sun et al., 2023) paper.
C.2 Hyperparameters
We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 4 and Table 5.
Table 4: Hyperparameters of SAMBO used in the D4RL datasets.
@@ C.2 Hyperparameters
walker2d-random 0.3 0.3 5 0
halfcheetah-medium 0.03 0.03 5 0
@@
hopper-medium-replay 0.05 0.05 5 0
walker2d-medium-replay 1 1 1 0
halfcheetah-medium-expert 0.3 0.3 5 0
hopper-medium-expert 1 1 5 0
walker2d-medium-expert 2 2 1 0
Table 5: Hyperparameters of SAMBO used in the NeoRL datasets.
Halfcheetah-L 0.01 0.01 5 0
Hopper-L 0.3 0.3 5 0
Walker2d-L 3 1 1 0
Halfcheetah-M 0.1 0.1 5 0
Hopper-M 0.3 0.3 5 0
Walker2d-M 3 1 1 0
Halfcheetah-H 5 1 5 0
Hopper-H 0.3 0.3 5 0
Walker2d-H 3 1 1 0
Coefficient $\alpha$ and $\beta$.
The model bias coefficient, $\alpha$, determines the degree of correction for model bias, while the policy shift coefficient, $\beta$, regulates the extent of correction for policy shift.
@@
We terminate training at 1M to prevent potential performance degradation due to the rapid convergence of the HalfCheetah environment in NeoRL. Additionally, to maintain training stability in the HalfCheetah-high dataset, we freeze classifiers after 0.5 million iterations. Similarly, in the walker2d-medium dataset of the D4RL benchmark, we ensure training stability by fixing the transition classifier after 1 million iterations.
C.3 Reward Function Discussion
Based on our derivation, if we start with $\mathbb{E}_{p^{\pi}(\tau)}e^{R(\tau)}$, we can obtain a new surrogate objective, which we refer to as SAMBO-r. In scenarios with sparse rewards and narrow datasets, where $\log r$ may distort the reward, SAMBO-r may be more suitable as a surrogate objective.
  • SAMBO. The modified reward is
    $\tilde{r}(s,a,s^{\prime})=\log r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
  • SAMBO-r. The modified reward is
    $\tilde{r}(s,a,s^{\prime})=r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
C.4 Computing Infrastructure
@@
Appendix D Omitted Experiments
D.1 Visualization
Figure 6: The $(s,a,s^{\prime})$ distributions generated by SAMBO result in the walker2d-medium-expert task in D4RL benchmark. Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
We also conducted a visualization experiment on the walker2d-medium-expert dataset within the D4RL benchmark. As shown in Figure 6, SAMBO effectively constrains the Model Data to closely match the Offline Data, and the Real Data generated by executing the learned policy in the actual environment does not significantly deviate from either the Offline Data or Model Data. This highlights the effectiveness of SAMBO in mitigating distribution shift across both benchmarks.
D.2 Detailed Ablation Discussion
SAMBO is regulated by the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
The model bias adjustment coefficient $\alpha$ controls the degree of adjustment to model-generated data. A larger $\alpha$ leads to excessive conservatism, restricting exploration even in well-performing regions and potentially causing suboptimal performance. Conversely, a smaller $\alpha$ may encourage excessive optimism in exploring risky regions, impairing performance. To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task. The results, shown in Fig. 7 (a), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance. Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration. This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy. A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy. However, this also introduces a heightened risk of taking unsafe actions. As shown in Fig. 7 (b), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness. A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability. Thus, $\beta$ should also be set appropriately to balance exploration and stability.
(a) Tuning for $\alpha$
(b) Tuning for $\beta$
Figure 7: Illustration of hyperparameter experiments in the HalfCheetah-random task in D4RL benchmark.
D.3 Performance Compare
We also compared the performance curves of model-based approaches on other tasks in the NeoRL benchmark. The curve for MOBILE was obtained by executing the official implementation. Since MOPO does not have an official implementation on NeoRL, we report its score based on the results presented in the original NeoRL paper.
As shown in Figures 10 to 16, SAMBO outperforms in all Hopper and Walker2d tasks, both of which are more complex and challenging. Although it exhibits slower convergence initially in the Walker2d-M and Walker2d-H tasks, its performance improves rapidly as the experiment progresses, ultimately achieving state-of-the-art (SOTA) levels. On the other hand, SAMBO also demonstrates comparable performance in simpler HalfCheetah tasks.
Figure 8: Hopper-L
Figure 9: Hopper-M
Figure 10: Hopper-H
Figure 11: Walker2d-L
Figure 12: Walker2d-M
Figure 13: Walker2d-H
Figure 14: HalfCheetah-L
Figure 15: HalfCheetah-M
Figure 16: HalfCheetah-H